40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17167     Entry time: Fri Sep 30 20:18:55 2022     In reply to: 17163     Reply to this: 17171
Author: Paco 
Type: Update 
Category: BHD 
Subject: LO phase noise with different actuation points 

[Paco, Koji]

We took lo phase noise spectra actuating on the for different optics-- LO1, LO2, AS1, and AS4. The servo was not changed during this time with a gain of 0.2, and we also took a noise spectrum without any light on the DCPDs. The plot is shown in Attachment #1, calibrated in rad/rtHz, and shown along with the rms values for the different suspension actuation points. The best one appears to be AS1 from this measurement, and all the optics seem to show the same 270 Hz (actually 268 Hz) resonant peak.


268 Hz noise investigation

Koji suspected the observed noise peak belongs to some servo oscillation, perhaps of mechanical origin so we first monitored the amplitude in an exponentially averaging spectrum. The noise didn't really seem to change too much, so we decided to try adding a bandstop filter around 268 Hz. After the filter was added in FM6, we turned it on and monitored the peak height as it began to fall slowly. We measured the half-decay time to be 264 seconds, which implies an oscillation with Q = 4.53 * f0 * tau ~ 3.2e5. This may or may not be mechanical, further investigation might be needed, but if it is mechanical it might explain why the peak persisted in Attachment #1 even when we change the actuation point; anyways we saw the peak drop ~ 20 dB after more than half an hour... After a while, we noticed the 536 Hz peak, its second harmonic, was persisting, even the third harmonic was visible.

So this may be LO1 violin mode & friends -

We should try and repeat this measurement after the oscillation has stopped, maybe looking at the spectra before we close the LO_PHASE control loop, then closing it carefully with our violin output filter on, and move on to other optics to see if they also show this noise.

Attachment 1: BHDC_asd_actuation_point.png  168 kB  Uploaded Fri Sep 30 21:42:37 2022  | Hide | Hide all
BHDC_asd_actuation_point.png
ELOG V3.1.3-